Steve traveled 600 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 1 hour less. Find the speed of his vehicle.

Answer by checkley77(12844)

D=RT
600=RT OR R=600/T
600=(R+20)(T-1) MULTIPLY THE 2 RIGHT TERMS.
600=RT+20T-R-20 REPLACE RT WITH 600 & R WITH 600/T
600=600+20T-600/T-20
0=20T-600/T-20
20T-600/T=20
(20T^2-600)/T=20
20T^2-600=20T
20T^2-20T-600=0
20(T^2-T-30)=0
20(T-6)(T+5)=0
T-6=0
T=6 HOURS FOR THE ORIGINAL TRIP.
600=R*6
600=6R
R=600/6
R=100 MPH.
PROOF:
600=(100+20)(6-1)
600=120*5
600=600